OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12

Parth (parthkohli):

The distance from zero of -2x + 8 is 12. That tells that -2x + 8 = \(\pm 12\) Two equations. \( \color{Black}{\Rightarrow -2x + 8 = -12 }\) \( \color{Black}{\Rightarrow -2x + 8 = 12 }\) Solve both.